OpenStudy (faiithhh):

PLEASE HELP ME SOLVE THIS INEQUALITY 8 |x + 3/4| < 2

OpenStudy (sshayer):

\[\left| x+a \right|<b,-b<x+a<b\]

OpenStudy (brainzonly):

Uh...I think use the distribution property first... 8x+6<2?

OpenStudy (brainzonly):

8x=-4 Now divide 8 on both sides. x=?

OpenStudy (faiithhh):

x would equal -.5 if that's the case, right?

OpenStudy (sshayer):

\[\left| x+\frac{ 3 }{ 4 } \right|<\frac{ 2 }{ 8 }\]

OpenStudy (sshayer):

\[\left| x+\frac{ 3 }{ 4 } \right|<\frac{ 1 }{ 4 }\] ?

OpenStudy (faiithhh):

im so lost tbh. i dont even know how to do this

OpenStudy (sshayer):

\[-\frac{ 1 }{ 4 }<x+\frac{ 3 }{ 4 }<\frac{ 1 }{ 4 }\] \[-\frac{ 1 }{ 4 } -\frac{ 3 }{ 4 }<x<\frac{ 1 }{ 4 }-\frac{ 3 }{ 4 }\] ?

OpenStudy (brainzonly):

umm...\[\frac{ -2 }{ 3 }\]

OpenStudy (brainzonly):

:D this is fun

OpenStudy (sshayer):

\[\frac{ -1-3 }{ 4 }<x<\frac{ 1-3 }{ 4 }\] \[\frac{ -4 }{ 4 }<x<\frac{ -2 }{ 4 }\] \[-1<x<\frac{ -1 }{ 2}\]

OpenStudy (sshayer):

or if there is positive constant ,then you can distribute \[\left| 8x+6 \right|<2\] -2<8x+6<2 -2-6<8x<2-6 -8<8x<-4 divide each inequality by 8 ?